# State-Constrained Offline Reinforcement Learning

### v1

@@ 6 Experimental results
Total (excl. rand) $432.4$ $602.2$ $677.4$ $692.6$ $628.3$ $697.7$ $753.1$ $\mathbf{813.4}$ 757.5
Antmaze Umaze $53.4$ $70.0$ $78.6$ $\mathbf{87.5}$ $25.0$ $\mathbf{88.6\pm 2.3}$ $75.6\pm 3.8$
U-diverse $64.6$ $44.0$ $\mathbf{71.4}$ $62.2$ $0.0$ $60.4\pm 26.4$ $67.2\pm 13.8$
M-play $0.0$ $0.0$ $3.0$ $\mathbf{71.2}$ $16.4$ $45.2\pm 26.0$ $20.6\pm 12.5$
M-diverse $0.8$ $0.0$ $10.6$ $\mathbf{70.0}$ $23.2$ $46.8\pm 26.2$ $13.0\pm 4.2$
L-play $0.0$ $0.0$ $0.0$ $\mathbf{39.6}$ $0.0$ $31.6\pm 5.0$ $5.2\pm 5.2$
L-diverse $0.0$ $0.0$ $0.2$ $\mathbf{47.5}$ $2.4$ $37.0\pm 21.5$ $2.2\pm 1.6$
Total (Antmaze) $118.8$ $114.0$ $163.8$ $\mathbf{378.0}$ $67.0$ $309.6$ $183.8$
Results for both StaCQ and our one-step version of StaCQ are shown in Table 1. Note that for the Antmaze tasks we do not need to train a reward model as Antmaze tasks are sparse rewards and so the reward for being in $s^{\prime}$ will remain the same. Despite the model error introduced in the state reachability metric and the policy by our simple models, StaCQ performs remarkably well against the baselines. StaCQ outperforms all the model-free baselines in the locomotion tasks (Hopper. Halfcheetah and Walker2d) on the whole range of datasets (random, medium-replay, medium and medium-expert). StaCQ also outperforms the model-based baselines on most of the locomotion tasks, except notably where RAMBO outperforms StaCQ in the Halfcheetah and random tasks. However on average across all locomotion tasks, StaCQ performs much better, where it is either the highest or second-highest performing method. Further, StaCQ outperforms RAMBO on the Antmaze tasks, which are notoriously difficult for methods that incorporate dynamics models [53, 42]; and compares well against the model-free baselines on these tasks. Importantly, StaCQ outperforms BCQ and TD3+BC which are the most comparable in theory and implementation. StaCQ performs consistently well in the Antmaze tasks, despite using dynamics models. The performance improvement of IQL over StaCQ in the Antmaze tasks is likely due to the advantage-weighted policy extraction method.
7 Discussion and conclusion
In this paper, we have introduced a method for state-constrained offline RL. The majority of prior offline RL approaches are batch-constrained, restricting the learning of the Q-function or the policy to $(s,a)$-pairs or the $(s,a)$-distribution in the dataset. In contrast, our state-constrained methodology confines learning updates solely to states within the dataset, offering significant potential reductions in dataset size for achieving optimality. By focusing on states rather than state-action pairs, the state-constrained approach allows for more efficient learning updates and can achieve optimality with smaller datasets compared to batch-constrained methods. Central to the proposed state-constrained approach is the concept of state reachability, which we rigorously define and exemplify with a simple maze environment. The maze example underscores the enhanced stitching capability of state-constrained methods over batch-constrained ones, especially with limited datasets that provide few trajectories. Further, we present theoretical backing affirming the potential for our method’s convergence, predicated on dataset states rather than state-action pairs. In addition, we have proposed StaCQ, an initial implementation of state-constrained deep offline RL. StaCQ leans on QS-learning [17] to estimate the value of transitioning from $s$ to $s^{\prime}$. This is a particularly advantageous approach in a state-constrained context as it avoids action dependency. StaCQ determines state reachability through forward and inverse dynamics models, refined via supervised learning on the dataset. While the performance of these models can influence StaCQ, the algorithm consistently demonstrates strong performance in complex environments, often surpassing baseline batch-constrained methods.
@@ 7 Discussion and conclusion
We envisage a number of potential improvements in future work. Despite the simplicity of the models we used, our reachability measure resulted in state-of-the-art performance. Future enhancements could further improve this performance by adding more complexity to the dynamics models, reducing model error, and capturing the true environment dynamics more effectively. Alternatively, developing entirely new state reachability criteria independent of dynamics models could significantly enhance the algorithm’s efficacy. Potential ideas include leveraging graph-based approaches or using reinforcement learning techniques to learn reachability directly from data. Moreover, the policy extraction process could be improved by incorporating elements such as a Decision Transformer [12] or diffusion policies [2]. These techniques have shown promising results in offline RL settings by leveraging the structure of the dataset and generating more diverse and effective policies. Integrating them into StaCQ could help the algorithm better exploit the available data and improve its overall performance. Similar to batch-constrained methods, further advances might arise from relaxed state constraints, albeit with tailored considerations for this setting. Utilising ensembles to assess the uncertainty of QA-values [3, 23, 7] and constraining actions to specific regions has proven advantageous in the batch-constrained domain. Similar methodologies could be extended to state-constrained approaches, such as employing ensembles of QS-functions to evaluate uncertainty.
@@
Beyond enhancing state reachability estimates, policy extraction mechanisms, and tweaking the state-constraint, the prospect for advancement in state-constrained methodologies is vast. One intriguing avenue is to refine state reachability estimates to exclusively focus on states, excluding action predictions. These advancements could give rise to algorithms that exclusively rely on state-only datasets. Such a paradigm is particularly relevant for real-world scenarios like video data where states are discernible (as images), but transitioning actions are unknown. Expanding the notion of state reachability is another promising frontier. Our current definition, rooted in one-step state reachability - $\exists a$ such that $p(s^{\prime}|s,a)=1$ can be extended to encompass multiple steps. For instance, a two-step reachability could be expressed as $s^{\prime\prime}\in\mathcal{SR}_{\mathcal{M}}(s)$ if $\exists a_{1},a_{2}$ such that $p(s^{\prime}|s,a_{1})=1$ and $p(s^{\prime\prime}|s^{\prime},a_{2})=1$. This principle can be further extrapolated, offering a foundation for inventive model-based offline RL algorithms. Lastly, the state-constrained methodology has significant potential implications for multi-agent offline RL scenarios, where the complexity grows with the size of the action space. By sidelining action dependence, our state-constrained approach seems well-suited for such challenges.
@@
To conclude, in this research, we have ventured into the realm of state-constrained offline reinforcement learning, introducing a novel paradigm that emphasises the importance of states over state-action pairs in the dataset. While this paper lays the foundational groundwork, the path forward calls for additional investigation, refinement, and possible integration with other emerging techniques. We believe that such endeavours can push the boundaries of current offline reinforcement learning methodologies and offer unprecedented solutions to complex problems.
@@ Appendix A Missing proofs
Which implies $|Q_{n+1}(s,s^{\prime})-Q^{*}(s,s^{\prime})|\leq\gamma\Delta_{n}$.
Where for brevity the maximisation $\max_{s^{\prime\prime}}$ is the shorthand for $\max_{s^{\prime\prime}\text{ s.t. }s^{\prime\prime}\in\mathcal{SR}_{\mathcal{M}}(s^{\prime})}$. Also, Eq. (15) is using the training rule; Eqs. (16) and (17) are simplifying and rearranging; (18) uses the fact that $|\max_{x}f_{1}(x)-\max_{x}f_{2}(x)|\leq\max_{x}|f_{1}(x)-f_{2}(x)|$; Eq. (19) we introduced a new variable $\hat{s}$ for which the maximisation is performed - this is permissible as allowing this additional variable to differ will always be at least the maximum value.
@@
In a deterministic setting, QA-values are equivalent to QS-values
@@
Case 1: In this situation, for all states in the dataset, we do not have any extra reachable states, other than the pairs i.e. $\forall s\in\mathcal{D},\mathcal{SR}_{\mathcal{M}}(s)=\{s^{\prime}\}$, where $(s,s^{\prime})\in\mathcal{D}$. In this case the state-constrained MDP, Definition 4.3, is equivalent to the batch-constrained MDP in [22]. The condition of the state-constrained MDP,
@@
Therefore, for case 1 BCQL and SCQL have the same Q-value update and therefore have the same optimal policy.
@@
This is the transition function for the batch-constrained MDP but allowing for an extra transition from $\tilde{s}$ to $\hat{s}^{\prime}$. Now comparing BCQL and SCQL Q-value updates, we have all equivalent values for the transition $(s,a,s^{\prime})$ except for the trajectory that contains the state $\tilde{s}$. For the converged QS-value, $Q^{*}$, let $\tilde{s}_{-1}$ be the state in the trajectory previous to $\tilde{s}$ and $\tilde{s}^{\prime}$ be the next state after $\tilde{s}$ in the trajectory,
@@
If $Q^{*}(\tilde{s},\tilde{s}^{\prime})\geq Q^{*}(\tilde{s},\hat{s}^{\prime})$ then again BCQL is equivalent to SCQL. However, if $Q^{*}(\tilde{s},\tilde{s}^{\prime})<Q^{*}(\tilde{s},\hat{s}^{\prime})$ then SCQL will have higher values than BCQL for all states previous to $\tilde{s}$ and therefore will produce a higher quality policy by the policy improvement theorem.
@@ Appendix B Reducing the complexity of state reachability estimation
For a given state $s$, we first determine its range by calculating the minimum and maximum values using $f_{\omega_{1}}$. Specifically, the minimum range $R_{\text{min}}(s)$ is computed as $\min_{i}f_{\omega_{1}}(s,a_{\text{rand}}^{i})$, and the maximum range $R_{\text{max}}(s)$ is $\max_{i}f_{\omega_{1}}(s,a_{\text{rand}}^{i})$. Subsequently, we construct a smaller set of states within the range $(R_{\text{min}}(s),R_{\text{max}}(s))$, using an R-tree [24], a data structure that efficiently identifies states within a specified hyper-rectangle (our range). This approach, leveraging the R-tree’s efficient search capability, dramatically reduces the size of the dataset. Consequently, the models are applied only to this refined set of states, leading to a significant reduction in computational complexity.

### v2

@@ 6 Experimental results
Total (excl. rand) $432.4$ $602.2$ $677.4$ $692.6$ $628.3$ $697.7$ $753.1$ $\mathbf{813.0}$ 757.5
Antmaze Umaze $53.4$ $70.0$ $78.6$ $\mathbf{87.5}$ $25.0$ $\mathbf{89.8\pm 5.0}$ $75.6\pm 3.8$
U-diverse $64.6$ $44.0$ $\mathbf{71.4}$ $62.2$ $0.0$ $64.0\pm 29.7$ $67.2\pm 13.8$
M-play $0.0$ $0.0$ $3.0$ $\mathbf{71.2}$ $16.4$ $47.2\pm 27.1$ $20.6\pm 12.5$
M-diverse $0.8$ $0.0$ $10.6$ $\mathbf{70.0}$ $23.2$ $47.4\pm 27.6$ $13.0\pm 4.2$
L-play $0.0$ $0.0$ $0.0$ $\mathbf{39.6}$ $0.0$ $31.4\pm 6.0$ $5.2\pm 5.2$
L-diverse $0.0$ $0.0$ $0.2$ $\mathbf{47.5}$ $2.4$ $40.0\pm 22.4$ $2.2\pm 1.6$
Total (Antmaze) $118.8$ $114.0$ $163.8$ $\mathbf{378.0}$ $67.0$ $319.8$ $183.8$
Results for both StaCQ and the one-step version of StaCQ are shown in Table 1. For the Antmaze tasks, it is unnecessary to train a reward model since these tasks feature sparse rewards, meaning the reward for being in $s^{\prime}$ remains constant. As all $s^{\prime}$ are in the dataset, the reward, $r(s,s^{\prime})$, is always known for all $s$ where $s^{\prime}\in\mathcal{SR}_{\mathcal{M}}(s)$. Also for the Antmaze tasks, StaCQ deploys independent target Q-value estimates. This is a techniques from Ghasemipour et al. (2022) that improves the performance of the policy in the medium and large Antmaze tasks. Despite potential model errors in estimating state reachability, StaCQ performs remarkably well against the baselines. The learning curves for all tasks in Table 1 are shown in Appendix G.
StaCQ outperforms all model-free baselines in the locomotion tasks (Hopper, Halfcheetah, and Walker2d) across the entire range of datasets (random, medium-replay, medium, and medium-expert). These environments are complex robotics tasks, with the random and medium-replay consisting of few good-quality episodes. StaCQ also surpasses most model-based baselines in the locomotion tasks, with the exception of RAMBO, which outperforms StaCQ in the Halfcheetah and random tasks. However, on average StaCQ is the highest-performing method across all locomotion tasks, consistently ranking as either the top or second-best performing method. TD3+BC uses a consistent hyperparameter across all datasets, whereas our StaCQ methods use an environment-dependent hyperparameter (see Appendix C). To enable a fair comparison, Appendix E compares TD3+BC and StaCQ using the same hyperparameter tuning strategy.
Through utilising the higher number of reachable states and thus finding higher quality OOD actions that lead to in-distribution states, StaCQ performs well across most of the locomotion tasks. However its performance in the Antmaze tasks is comparatively lower. StaCQ does outperform RAMBO in the Antmaze tasks, which are notoriously challenging for methods involving dynamics models (Wang et al., 2021; Rigter et al., 2022), and performs competitively against the model-free baselines. Notably, StaCQ outperforms BCQ and TD3+BC, which are the most comparable baselines in terms of both theory and implementation. Methods such as IQL, which rely on advantage-weighted policy extraction, may have an advantage in these tasks due to their ability to handle sparse rewards more effectively. Nonetheless, StaCQ is either the highest or second highest performing method on the Antmaze tasks against the comparative methods.
7 Discussion and conclusions
In this paper, we introduced a method for state-constrained offline RL. Most prior offline RL approaches are batch-constrained, limiting the learning of the Q-function or policy to $(s,a)$ pairs or the $(s,a)$-distribution in the dataset. In contrast, our state-constrained methodology confines learning updates to states within the dataset, offering significant potential reductions in dataset size required for achieving optimality. By focusing on states rather than state-action pairs, the state-constrained approach enables more efficient learning updates and can achieve optimality with smaller datasets compared to batch-constrained methods.
Central to the proposed state-constrained approach is the concept of state reachability, which we define and illustrate in a simple maze environment. This example underscores the improved stitching capability of state-constrained methods over batch-constrained ones, particularly in datasets with limited trajectories. Additionally, we provide theoretical backing that affirms the potential for our method’s convergence based on dataset states rather than state-action pairs. Similar to BCQ (Fujimoto et al., 2019), our theory assumes a deterministic MDP. This focus on deterministic environments simplifies the theoretical development (Fujimoto et al., 2019; Jin et al., 2021; Shi et al., 2022), and provides a fair comparison with the BCQ theory, while also providing clear guarantees for convergence, making it an appropriate starting point for the state-constrained approach. Our theoretical goal is to demonstrate convergence and provide a comparison with BCQ. As a result, Theorem 3.6 shows that SCQL converges to the optimal policy with a smaller requisite dataset size than BCQ; and, Theorem 3.7 shows that the policy produced by SCQL can always be preferred to the policy produced by BCQL.
Based from these theoretical developments, our practical implementation also defines reachable states deterministically — states are considered reachable or not, regardless of potential stochastic transitions. Deterministic settings align with many practical applications, such as robotic control or industrial systems, where transitions can be accurately modelled. However, some real-world environments are stochastic, where actions lead to a distribution of possible next states. Future work will extend state reachability to account for stochastic transitions by incorporating probabilistic models of the environment’s dynamics. Theoretically, the definition of state reachability can be extended to stochastic MDPs, for example $s^{\prime}$ is reachable from $s$ if it is the most likely next state when performing action $a$, i.e $s^{\prime}\in\mathcal{SR}_{\mathcal{M}}(s)\text{ if }p(s^{\prime}|s,a)=\max_{s^{\prime}_{i}\sim p(\cdot|s,a)}p(s^{\prime}_{i}|s,a)$. Practically, this will involve redefining reachability in terms of the likelihood of reaching a state under certain actions, possibly using techniques such as probabilistic graphical models. The key challenge will be maintaining computational efficiency while handling the increased complexity of probabilistic transitions. Extending the framework in this way will allow the state-constrained approach to be applied to more complex, uncertain environments, improving its robustness and real-world applicability.
We also introduced StaCQ, an implementation of state-constrained deep offline RL. StaCQ leverages QSS-learning (Edwards et al., 2020) to estimate the value of transitioning from $s$ to $s^{\prime}$, which is particularly advantageous in a state-constrained context as it avoids action dependency. StaCQ determines state reachability using forward and inverse dynamics models, trained via supervised learning on the dataset. While the performance of these models can influence StaCQ, the algorithm consistently demonstrates strong performance in complex environments, often surpassing baseline batch-constrained methods. StaCQ relies on an ensemble of forward and inverse dynamics models. Increasing the size of the ensemble would enable more accurate estimates, however the number of models is consistent with prior research (Janner et al., 2019) and is accurate enough for StaCQ to perform well, as shown in Table 1. StaCQ uses an ensemble of four critics whereas OneStep StacQ uses a single critic. These values were chosen as they enable stable estimates across all datasets.
@@ 7 Discussion and conclusions
There are several potential improvements to explore in future work. Despite the simplicity of the models we used, our reachability measure delivered SOTA performance. Future enhancements could improve this by adding more complexity to the dynamics models, reducing model error, and capturing the true environment dynamics more accurately. Alternatively, developing entirely new state reachability criteria independent of dynamics models could significantly enhance the algorithm’s effectiveness. Potential ideas include leveraging graph-based approaches or using reinforcement learning techniques to learn reachability directly from data. Additionally, the policy extraction process could be improved by incorporating elements such as a Decision Transformer (Chen et al., 2021) or diffusion policies (Ajay et al., 2022). These techniques have shown promise in offline RL by leveraging the structure of the dataset to generate more diverse and effective policies. Integrating them into StaCQ could allow the algorithm to exploit the available data more effectively.
Similar to batch-constrained methods, further advances might arise from relaxed state constraints, albeit with tailored considerations for this setting. Using ensembles to assess the uncertainty of QSA-values (An et al., 2021; Ghasemipour et al., 2022; Beeson & Montana, 2024) and constraining actions to specific regions has proven advantageous in the batch-constrained domain. Similar methodologies could be extended to state-constrained approaches, such as employing ensembles of QSS-functions to evaluate uncertainty.
@@
Beyond improving state reachability estimates and policy extraction mechanisms, the potential for advancement in state-constrained methodologies is vast. One intriguing direction is to refine state reachability estimates to focus solely on states, excluding action predictions. Such advancements could lead to algorithms that rely entirely on state-only datasets, which is particularly relevant for real-world scenarios like learning from video data, where states are observable (as images), but the actions leading to transitions are unknown. Expanding the notion of state reachability is another promising avenue. Our current definition, based on one-step state reachability ($\exists a$ such that $p(s^{\prime}|s,a)=1$), could be extended to multiple steps. For example, a two-step reachability could be defined as $s^{\prime\prime}\in\mathcal{SR}_{\mathcal{M}}(s)$ if $\exists a_{1},a_{2}$ such that $p(s^{\prime}|s,a_{1})=1$ and $p(s^{\prime\prime}|s^{\prime},a_{2})=1$. This principle can be further extrapolated, providing a foundation for novel model-based offline RL algorithms. Extending this notion of state reachability will allow for a higher number of reachable states to be found, while being sure they lead, in $k$-steps, to in-distribution states. Appendix H shows the theoretical benefits of a multi-step state reachable method on a simple maze environment. However, a practical method that incorporates model error is left for future work. Lastly, the state-constrained methodology holds significant potential for multi-agent offline RL scenarios, where complexity grows with the action space size. By reducing action dependence, the state-constrained approach appears well-suited to tackle such challenges.
@@
In conclusion, we have ventured into the realm of state-constrained offline RL, introducing a novel paradigm that emphasises states over state-action pairs in the dataset. While this paper lays the foundational groundwork, the path forward calls for further investigation, refinement, and integration with other emerging techniques. We believe that these endeavours can push the boundaries of current offline RL methodologies and offer unprecedented solutions to complex problems.
@@ Appendix A Missing proofs
Where for brevity the maximisation $\max_{s^{\prime\prime}}$ is the shorthand for $\max_{s^{\prime\prime}\text{ s.t. }s^{\prime\prime}\in\mathcal{SR}_{\mathcal{M}}(s^{\prime})}$. Also, Eq. equation 16 is using the training rule; Eqs. equation 17 and equation 18 are simplifying and rearranging; equation 19 uses the fact that $|\max_{x}f_{1}(x)-\max_{x}f_{2}(x)|\leq\max_{x}|f_{1}(x)-f_{2}(x)|$; Eq. equation 20 we introduced a new variable $\hat{s}$ for which the maximisation is performed - this is permissible as allowing this additional variable to differ will always be at least the maximum value.
Thus, the updated $Q_{n+1}(s,s^{\prime})$ for any $s,s^{\prime}$ is at most $\gamma$ times the maximum error in the $Q_{n}$ table, $\Delta_{n}$. The largest error in the initial table, $\Delta_{0}$, is bounded because the values of $Q_{0}(s,s^{\prime})$ and $Q^{*}(s,s^{\prime})$ are bounded $\forall s,s^{\prime}.$ Now, after the first interval during which each $s,s^{\prime}$ is visited the largest error will be at most $\gamma\Delta_{0}$. After $k$ such intervals, the error will be at most $\gamma^{k}\Delta_{0}$. Since each state is visited infinitely often, the number of such intervals is infinite and $\Delta_{n}\rightarrow 0$ as $n\rightarrow\infty$.
@@
In a deterministic setting, QSA-values are equivalent to QSS-values
@@
Case 1: In this situation, for all states in the dataset, we do not have any extra reachable states, other than the pairs i.e. $\forall s\in\mathcal{D},\mathcal{SR}_{\mathcal{M}}(s)=\{s^{\prime}\}$, where $(s,s^{\prime})\in\mathcal{D}$. In this case the state-constrained MDP, Definition 3.3, is equivalent to the batch-constrained MDP in (Fujimoto et al., 2019). The condition of the state-constrained MDP,
@@
Therefore, for case 1 BCQL and SCQL have the same Q-value update and therefore have the same optimal policy. So, $\pi_{\text{BCQL}}(s)=\pi_{\text{SCQL}}(s),Q(s,\pi_{\text{BCQL}}(s))=Q(s,\pi_{\text{SCQL}}(s)),\text{and thus }V_{\pi_{\text{BCQL}}}(s)=V_{\pi_{\text{SCQL}}}(s),\forall s\in\mathcal{D}$.
@@
This is the transition function for the batch-constrained MDP but allowing for an extra transition from $\tilde{s}$ to $\hat{s}^{\prime}$. Now comparing BCQL and SCQL Q-value updates, we have all equivalent values for the transition $(s,a,s^{\prime})$ except for the trajectory that contains the state $\tilde{s}$. For the converged QSS-value, $Q^{*}$, let $\tilde{s}_{-1}$ be the state in the trajectory previous to $\tilde{s}$ and $\tilde{s}^{\prime}$ be the next state after $\tilde{s}$ in the trajectory,
@@
If $Q^{*}(\tilde{s},\tilde{s}^{\prime})\geq Q^{*}(\tilde{s},\hat{s}^{\prime})$ then again BCQL is equivalent to SCQL. However, if $Q^{*}(\tilde{s},\tilde{s}^{\prime})<Q^{*}(\tilde{s},\hat{s}^{\prime})$ then SCQL will have higher values than BCQL for all states previous to $\tilde{s}$ and therefore will produce a higher quality policy by the policy improvement theorem. In this case, as $Q^{*}(\tilde{s},\tilde{s}^{\prime})<Q^{*}(\tilde{s},\hat{s}^{\prime})$, $\exists s$ such that $Q(s,\pi_{\text{SCQL}}(s))>Q(s,\pi_{\text{BCQL}}(s))$, thus $V_{\pi_{\text{SCQL}}}(s)\geq V_{\pi_{\text{BCQL}}}(s),\forall s\in\mathcal{D}$.
@@ Appendix B Reducing the complexity of state reachability estimation
For a given state $s$, we first determine its range by calculating the minimum and maximum values using $f_{\omega_{1}}$. Specifically, the minimum range $R_{\text{min}}(s)$ is computed as $\min_{i}f_{\omega_{1}}(s,a_{\text{rand}}^{i})$, and the maximum range $R_{\text{max}}(s)$ is $\max_{i}f_{\omega_{1}}(s,a_{\text{rand}}^{i})$. Subsequently, we construct a smaller set of states within the range $(R_{\text{min}}(s),R_{\text{max}}(s))$, using an R-tree (Guttman, 1984), a data structure that efficiently identifies states within a specified hyper-rectangle (our range). This approach, leveraging the R-tree’s efficient search capability, dramatically reduces the size of the dataset. Consequently, the models are applied only to this refined set of states, leading to a significant reduction in computational complexity.